OpenStudy (anonymous):

I need help in Analytical Geometry Find the equation of a circle tangent to the lines 2x+y=4, 2x+y=2 and x-2y+5=0

OpenStudy (yamyam70):

@Nnesha can you help my friend?

OpenStudy (yamyam70):

@nincompoop

OpenStudy (yamyam70):

@Lord2clash @ConnexusStudent0912 @chainedecho can someone help with the geometry?

OpenStudy (phi):

Have you done any similar problems to this one ? I can guess at an approach, but it seems complicated.

OpenStudy (anonymous):

not yet. our professor gave this as our assignment

OpenStudy (phi):

I would plot the 3 lines, and they should make a triangle. Now the problem is finding the "in-circle" of a triangle.

OpenStudy (phi):

the center of the circle is where the "angle bisectors" (of the triangle) meet.

OpenStudy (yamyam70):



OpenStudy (phi):

though looking at the equations 2x+y=4, 2x+y=2 x-2y+5=0 the first 2 are parallel, so the distance between those two lines is the diameter of the circle

OpenStudy (yamyam70):

@phi is my figure right?

OpenStudy (anonymous):



OpenStudy (anonymous):

is this correct?

OpenStudy (phi):

yes, here is a geogebra plot (geogebra is a free download)

OpenStudy (phi):

the center of the circle will be on the line that bisects the two parallel lines it will be a distance of 1 radius from the perpendicular line. because we can go either "up" or "down", there are two different points that could be the center of the circle.

OpenStudy (anonymous):

can you help me with the solution to find the general equation of that problem?

OpenStudy (phi):

The first step is to find the distance between the two parallel lines any idea how to do that ? we could use vectors if you have learned that approach. or, if you look at your 3 lines in slope-intercept form: y= -2x +4 y= -2x +2 y= (½) x + 5/2 you see the 3rd line has a slope of ½ which is the negative reciprocal of -2 (the slope of the two other lines) so we could find the intersection points and then find the distance between two points using the distance formula.

OpenStudy (phi):

* the 3rd line is perpendicular to the other two lines

OpenStudy (anonymous):

\[\sqrt{(x _{2}-x _{1)}+(y _{2}-y _{1)}}\]

OpenStudy (anonymous):

is this the distance formula that you're talking about or the one with r^2?

OpenStudy (phi):

yes, that is the distance formula but first we need the two points

OpenStudy (phi):

you can find point A by solving 2x+y=4 x-2y= -5 multiply the 2nd equation by -2 (all terms, both sides) and you get -2x+4y= 10 add the two equations term by term 2x+y=4 -2x+4y= 10 ----------- 0 + 5y= 14 y=14/5 or 2.8 and using the 2nd equation, x -2*2.8= -5 x= -5+5.6 = 0.6 so point A is at (0.6, 2.8)